Mr. D.B. Carpenter Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S018 1GS

May 4, 1994

TV Licensing, Bristol, BS98 1TL

Dear TV Licensing,

I have received another TV Licence Enquiry (reference 1157832635) claiming that you had no reply to the previous one. I certainly did reply to it (promptly). The reply contained notification of my changed address. At the time of the previous enquiry I still had a valid licence quoting my old address

Mr. D.B. Carpenter 98 Gordon Avenue Southampton SO2 1DA

That licence expired on the 30th April. If you will remove my old address from your records (to stop me getting more forwarded reminders) and send me a new licence dated from the first of this month, I will be happy to pay the fee.

Yours sincerely,